# Optimal Control of Behavioral-Feedback SIR Epidemic Model

### v1

@@ 5 Counterexample
To illustrate this, we compare the cost of the filling-the-box strategy for two different thresholds $\bar{y}$. The initial condition is fixed at $y_{0}=0.08$ and $x_{0}=1-y_{0}$. Figure 3 shows the two simulations. In each case, the top plot displays the evolution of the infection curve: the uncontrolled trajectory is represented by a blue dashed line, while the controlled trajectory obtained with the filling-the-box strategy is plotted in green. The bottom plot shows the corresponding control function applied over time.
When the threshold is set to $\bar{y}=0.11$, the policy produces a total cost of $47.7$. Surprisingly, increasing the threshold to $\bar{y}=0.154$ leads to a higher total cost of $51.44$. As the control signal implemented for $\bar{y}=0.11$ is also feasible for $\bar{y}=0.154$, this proves that the filling-the-box strategy is not optimal for this transmission rate. The explanation for this fact is that, since the reproduction number in this example is increasing in $y$, keeping the infection curve at the highest admissible level allows to deactivate the control sooner, but requires a stronger intervention (as the bottom plots of Figure 3 illustrate), resulting in a larger cumulated cost. This example shows that when Assumption 1 is not satisfied, the simple filling-the-box strategy may fail to be optimal. In such situations, the timing of interventions becomes critical, and optimal policies may require more sophisticated strategies than a fixed-threshold rule.
@@ Appendix A Proof of Proposition 1
The right-hand side of (1) is locally Lipschitz with respect to $(x,y,z)$. Hence, on every time interval on which $u$ is continuous, local existence and uniqueness of the solution for the corresponding Cauchy problem is a consequence of the Picard-Lindelöf theorem [hale, I.3]). Consider now any interval $\mathcal{I}$ admitting minimum $t_{0}$ on which $u$ is continuous. By integrating equations in (1), assuming, for $t=t_{0}$, initial condition $(x_{0},y_{0},z_{0})$ in $\Delta$, we get
@@
In other terms, as long as the solution exists, it lives in $\Delta$. In particular, this implies that the solution is globally defined on the whole of $\mathcal{I}$ [knauf, Section 3.3]. We now construct a global solution on $0,+\infty)$ in a recursive way. We order the discontinuity points as $t_{1}<t_{2}<\cdots$. First, we let $u^{k}:[t_{k},t_{k+1}]\to\mathbb{R}$ for $k\geq 0$ (letting $t_{0}=0$) to be the unique continuous function that coincides with $u$ on $t_{k},t_{k+1}]$. Similarly, we define $u^{k}:[t_{k},+\infty)\to\mathbb{R}$ when the set of discontinuities is finite and has cardinality $k$. We now define $(x(t),y(t),z(t))$ on $t_{0},t_{1}]$ as the unique solution of the Cauchy problem relative to (1) with $u=u_{0}$ and initial condition $(x_{0},y_{0},z_{0})$. Assuming the solution $(x,y,z)$ to be constructed up to $t_{k}$ with the desired properties, we extend it to the next discontinuity (or to $\infty$ in case this was the last one) by considering the solution $(x^{k}(t),y^{k}(t),z^{k}(t))$ of the Cauchy problem on $t_{k},t_{k+1}]$ relative to the input function $u_{k}$ and initial condition $(x(t_{k}),y(t_{k}),z(t_{k}))$. We then extend $(x,y,z)$ to $0,t_{k+1}]$ by setting $(x(t),y(t),z(t))=(x^{k}(t),y^{k}(t),z^{k}(t))$ for $t\in[t_{k},t_{k+1}]$. This construction yields the first part of the proposition.
@@
[[Uncaptioned image]] Francesca Parise received the B.Sc. and M.Sc. (summa cum laude) degrees in information and automation engineering from the University of Padova, Padua, Italy, in 2010 and 2012, respectively, the graduate degree from the Galilean School of Excellence, University of Padova, in 2013, and the Ph.D. degree in information technology and electrical engineering from the Automatic Control Laboratory, ETH Zurich, Zurich, Switzerland, in 2016. She conducted her master’s thesis research with Imperial College London, London, U.K., in 2012. From 2016 to 2020, she was a Postdoctoral Researcher with the Laboratory for Information and Decision Systems, MIT, and is currently an Assistant Professor of Electrical and Computer Engineering with Cornell University, Ithaca, NY. Her research interests include identification, analysis, and control of multiagent systems, with application to transportation, social, economic networks, and systems biology.

### v2

@@ 5 Counterexample
To illustrate this, we compare the cost of the filling-the-box strategy for two different thresholds $\bar{y}$. The initial condition is fixed at $y_{0}=0.08$ and $x_{0}=1-y_{0}$. Figure 3 shows the two simulations.
When the threshold is set to $\bar{y}=0.11$, the policy produces a total cost of $47.7$. Surprisingly, increasing the threshold to $\bar{y}=0.154$ leads to a higher total cost of $51.44$. As the control signal implemented for $\bar{y}=0.11$ is also feasible for $\bar{y}=0.154$, this proves that the filling-the-box strategy is not optimal for this transmission rate. The explanation for this fact is that, since the reproduction number in this example is increasing in $y$, keeping the infection curve at the highest admissible level does still lead to shorter infection (i.e. one can deactivate the control sooner), but it requires a stronger intervention (as the bottom plots of Figure 3 illustrate), resulting in a larger cumulated cost. This example shows that when Assumption 1 is not satisfied, the simple filling-the-box strategy may fail to be optimal, so that no further extension of our optimality result is possible in general. At the same time, we have identified a broad and realistic class of infection-rate functions for which Assumption 1 is satisfied, implying that the filling-the-box policy remains optimal in many practically relevant scenarios, while outside this regime the timing of interventions becomes critical and optimal policies may require more sophisticated strategies than a fixed-threshold rule.
@@ Appendix A Proof of Proposition 1
The right-hand side of (1) is locally Lipschitz with respect to $(x,y,z)$. Hence, on every time interval on which $u$ is continuous, local existence and uniqueness of the solution for the corresponding Cauchy problem is a consequence of the Picard-Lindelöf theorem [20, I.3]). Consider now any interval $\mathcal{I}$ admitting minimum $t_{0}$ on which $u$ is continuous. By integrating equations in (1), assuming, for $t=t_{0}$, initial condition $(x_{0},y_{0},z_{0})$ in $\Delta$, we get
@@
In other terms, as long as the solution exists, it lives in $\Delta$. In particular, this implies that the solution is globally defined on the whole of $\mathcal{I}$ [27, Section 3.3]. We now construct a global solution on $0,+\infty)$ in a recursive way. We order the discontinuity points as $t_{1}<t_{2}<\cdots$. First, we let $u^{k}:[t_{k},t_{k+1}]\to\mathbb{R}$ for $k\geq 0$ (letting $t_{0}=0$) to be the unique continuous function that coincides with $u$ on $t_{k},t_{k+1}]$. Similarly, we define $u^{k}:[t_{k},+\infty)\to\mathbb{R}$ when the set of discontinuities is finite and has cardinality $k$. We now define $(x(t),y(t),z(t))$ on $t_{0},t_{1}]$ as the unique solution of the Cauchy problem relative to (1) with $u=u_{0}$ and initial condition $(x_{0},y_{0},z_{0})$. Assuming the solution $(x,y,z)$ to be constructed up to $t_{k}$ with the desired properties, we extend it to the next discontinuity (or to $\infty$ in case this was the last one) by considering the solution $(x^{k}(t),y^{k}(t),z^{k}(t))$ of the Cauchy problem on $t_{k},t_{k+1}]$ relative to the input function $u_{k}$ and initial condition $(x(t_{k}),y(t_{k}),z(t_{k}))$. We then extend $(x,y,z)$ to $0,t_{k+1}]$ by setting $(x(t),y(t),z(t))=(x^{k}(t),y^{k}(t),z^{k}(t))$ for $t\in[t_{k},t_{k+1}]$. This construction yields the first part of the proposition.
@@
[[Uncaptioned image]] Francesca Parise is an Assistant Professor at Cornell University in the School of Electrical and Computer Engineering. She received the B.Sc. and M.Sc. degrees in Information and Automation Engineering from the University of Padova, Italy, in 2010 and 2012. She graduated from the Galilean School of Higher Education, University of Padova, Italy, in 2013. She defended her Ph.D. at the Automatic Control Laboratory, ETH Zurich, Switzerland in 2016 and was a Postdoctoral researcher at MIT from 2016 to 2020. Her research focuses on analysis and control of large multi-agent systems, with application to transportation, energy, social and economic networks.